OpenStudy (anonymous):

subtract: t^2+7t/t^2+8t+7 subtracted from 5t+35/t^2+8t+7

jhonyy9 (jhonyy9):

t2 +7t=t(t+7) t2 +8t+7=(t+1)(t+7) 5t+35=5(t+5) t(t+7) 5(t+7) ---------- - ----------- = t/(t+1) - 5/(t+1) = (t-5)/(t+1) (t+1)(t+7) (t+1)(t+7)